OpenStudy (anonymous):

HELPPPPPPPPPPPP Part A: Solve F = 9 over 5C + 32 for C. (4 points) Part B: Determine the value of C when F = 10 °F. (2 points) Part C: Solve –np – 70 > 40 for n. Show your work.

OpenStudy (anonymous):

WILL GIVE MEDAL

OpenStudy (anonymous):

anybody please

OpenStudy (anonymous):

@tsade500

OpenStudy (anonymous):

a) F = (9/5)C + 32 F - 32 = (9/5)C (5/9)(F - 32) = C b) -12.22 c) –np – 70 > 40 -np>110 n< -110/p

OpenStudy (anonymous):

thanks